OpenStudy (anonymous):

how do I integrate x^3(sin5x)

OpenStudy (tkhunny):

Parts, multiple times. \(\int x^{3}\sin(5x)\;dx = \int x^{3}\;d\left[-\dfrac{1}{5}\cos(5x)\right]\) Keep going!

OpenStudy (kirbykirby):

I think it's worth noting about "tabular integration by parts" which summarizes multiple integration by parts more rapidly: http://www.kkuniyuk.com/Math151TurboParts.pdf

OpenStudy (akashdeepdeb):

Try integrating by parts, Twice! Integration by parts says: \[\int u.dv = uv - \int v.du\] So, to solve this question: \[\int x^3~sin~5x~dx = -x^3~cos~ 5x + \int 3x^2~cos~5x~dx\] You get that ^ when you set \(u = x^3 \) and \(dv = sin ~5x~dx\) NOW We use the second term on the right hand side. ie. \(\int 3x^2~cos~5x~dx\) Now integrate this by parts by setting \(u = cos ~5x\) \(dv = 3x^2.dx\) Calculate this integral, and then substitute it in your original formula. You should get the answer. Did you get this? :)

OpenStudy (anonymous):

uhhh none of these answers are really making sense to me, is there a table equation im supposed to be using or something?

OpenStudy (tkhunny):

Have you done any integration by parts?

OpenStudy (anonymous):

omg its been a while but i just remembered the formula you have to plug it into. haha lemme try again

OpenStudy (tkhunny):

While you are catching up, you may wish to review the material provided by kirbykirby. It requires Way - WAY less algebra than the version I presented. And my version is FAR less algebra than the crazy, laborious, American way AkashdeepDeb suggested.

OpenStudy (kirbykirby):

^Hehe true. I only discovered integration by parts by myself though sort of randomly, after I had done all my calculus courses . What a drag -_-

OpenStudy (kirbykirby):

^well the tabular method is what I meant hehe!

OpenStudy (akashdeepdeb):

Wow! That method is great! Thanks. :)

OpenStudy (anonymous):

wow that method seriously is great, thank you guys so much!